OpenStudy (anonymous):

Will give medal. How do I get -x + 5y =3 into slope-intercept form? Show work please so I understand.

OpenStudy (dmndlife24):

add x on both sides to get the y on the left side... y = mx + b m = slope and b = y int

OpenStudy (dmndlife24):

divide by 5 to isolate y... y = 1/5x + 3/5

OpenStudy (anonymous):

\[-x + x +5y = 3 + x \]

OpenStudy (dmndlife24):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (dmndlife24):

np

OpenStudy (anonymous):

slope intercept form is y=mx+b. you want to get y by itself so -x+5y=3 add x to both sides 5y=x+3 divide both sides by 5 you get y = 1/5x + 3/5 im very sure this is correct, tell me if im wrong :)